OpenStudy (anonymous):

Which property allows you to state if m<'4 + <'5 = 59°, and m<4 = 20°, then 20° + m<'5 = 59 @vzfreakz

OpenStudy (anonymous):

substitution

OpenStudy (anonymous):

i think?

OpenStudy (anonymous):

Addition and Substitution Property.

OpenStudy (anonymous):

Wait, it's either Substitution, or Addition Property.

OpenStudy (anonymous):

@vzfreakz it gives me (addition and subtraction properties) or substitution property?

OpenStudy (anonymous):

@vzfreakz those are my awnser choices

OpenStudy (anonymous):

im thinking only substitution, because you're not adding anything.

OpenStudy (anonymous):

@gagafacea1 if m<1 +m<2=90 and m<2=m<3 then m<1+m<3=90 what property is it?

OpenStudy (anonymous):

again, it's substitution!

OpenStudy (anonymous):

Let's go with substitution.

OpenStudy (anonymous):

@gagafacea1 if 1/2AB= 1/2CD which property can be used to justify the statement AB = CD?